Suppose an economy is faced with a $20 billion dollar recessionary gap. Furthermore, the marginal propensity to consume is .80 in the economy. If taxes are decreased by $1 billion dollars, then government spending needs to increase by _____________ billion?
1
Expert's answer
2020-12-03T07:01:51-0500

Solution:

MPS = 0.80


Derive MPC:

"MPC+MPS=1"

"MPC =1-MPS"

"MPC=1-0.80"

"MPC=0.20"


Calculate the spending multiplier:

"Multiplier=\frac{1}{MPS}"


"Multiplier = \frac{1}{0.80} = 1.25"


"Increase\; in \;government \;spending=\frac{recessionary gap}{spending multiplier}"


="\frac{20}{1.25} = 16 \;billion dollars"



Increase in Government spending needed = 16 Billion dollars